# A Unified Framework for Forward and Inverse Problems in Subsurface Imaging using Latent Space Translations

### v1

@@ Abstract
In subsurface imaging, learning the mapping from velocity maps to seismic waveforms (forward problem) and waveforms to velocity (inverse problem) is important for several applications. While traditional techniques for solving forward and inverse problems are computationally prohibitive, there is a growing interest in leveraging recent advances in deep learning to learn the mapping between velocity maps and seismic waveform images directly from data. Despite the variety of architectures explored in previous works, several open questions still remain unanswered such as the effect of latent space sizes, the importance of manifold learning, the complexity of translation models, and the value of jointly solving forward and inverse problems. We propose a unified framework to systematically characterize prior research in this area termed the Generalized Forward-Inverse (GFI) framework, building on the assumption of manifolds and latent space translations. We show that GFI encompasses previous works in deep learning for subsurface imaging, which can be viewed as specific instantiations of GFI. We also propose two new model architectures within the framework of GFI: Latent U-Net and Invertible X-Net, leveraging the power of U-Nets for domain translation and the ability of IU-Nets to simultaneously learn forward and inverse translations, respectively. We show that our proposed models achieve state-of-the-art (SOTA) performance for forward and inverse problems on a wide range of synthetic datasets, and also investigate their zero-shot effectiveness on two real-world-like datasets.
@@ 1 Introduction
The goal of subsurface imaging is to identify the structure and geophysical properties of layers underneath the Earth’s surface such as velocity maps ($v$), which is useful for a number of applications including energy exploration, carbon capture and sequestration, and developing earthquake early warning systems (Zhang et al., 2013). A typical approach for subsurface imaging is to conduct seismic surveys, where the elastic wave energy generated from one or more controlled sources is made to propagate through the Earth’s layers that are refracted, reflected, and received as seismic waveforms ($p$) on the surface by multiple receivers. Mathematically, the relation between velocity maps and seismic waveforms is governed by the following acoustic wave equation expressed as a non-linear partial differential equation (PDE):
@@
For example, several deep learning methods have recently been proposed for FWI using encoder-decoder architectures such as InversionNet (Wu & Lin, 2019) and VelocityGAN (Zhang et al., 2019), where the encoder maps transform seismic data to latent feature representations that are decoded back to the velocity domain (Lin et al., 2023). Techniques for solving the forward problem in PDEs include the seminal work on physics-informed neural networks (PINNs) (Raissi et al., 2019) and recent works in neural operators such as Fourier Neural Operators (FNOs) (Li et al., 2020). The basic idea of FNOs is to lift inputs to the frequency domain using Fourier transform and apply linear transforms before inverting them to the original space in every FNO layer. A recent work has also pursued the goal of solving both forward and inverse problems together using a common architecture termed Auto-Linear (Feng et al., 2024), where two auto-encoders are first separately trained to project velocity and waveform fields to higher-dimensional manifolds, which are then linked using linear transforms for translating between manifolds.
@@ 2.1 Deep Learning for Subsurface Imaging (DL4SI)
Since ground-truth observations of velocity maps are difficult to obtain in real-world settings, a number of unsupervised and self-supervised formulations have also been explored for FWI. This includes the framework of UPFWI (Jin et al., 2021) that uses the supervision of a physics-based forward model of the wave equation to enforce cycle-consistency with a data-driven inverse model. While the training of UPFWI is completely unsupervised and achieves comparable performance as InversionNet in some settings, it requires computationally expensive simulations of a forward model at every training epoch, restricting its scalability. A recent line of work in semi-supervised learning for FWI involves the framework of Auto-Linear (Feng et al., 2024). In this framework, both velocity and waveforms are first mapped to higher-dimensional manifolds using two auto-encoders separately trained to minimize reconstruction loss in their respective domains. This is followed by a supervised learning of translations between the manifolds using two independent linear transforms (one for forward and one for inverse), building on previous observations of linearity in high-dimensional spaces for FWI (Feng et al., 2022).
@@ 3.1 Generalized Forward Inverse Framework
As shown schematically in Figure 1, our proposed framework of GFI unifies the learning of the forward problem $f_{v\rightarrow p}:\mathcal{V}\rightarrow\mathcal{P}$ and the inverse problem $f_{p\rightarrow v}:\mathcal{P}\rightarrow\mathcal{V}$ using a common architecture involving latent space translations.
The GFI framework is inspired by two key assumptions. First, according to the manifold assumption, we assume that the velocity maps $v\in\mathcal{V}$ and seismic waveforms $p\in\mathcal{P}$ can be projected to their corresponding latent space representations, $\tilde{v}$ and $\tilde{p}$, respectively, which can be mapped back to their reconstructions in the original space, $\hat{v}$ and $\hat{p}$. Note that the sizes of the latent spaces can be smaller or larger than the original spaces. Further, the size of $\tilde{v}$ may not match with the size of $\tilde{p}$ (e.g., in Auto-Linear). Second, according to the latent space translation assumption, we assume that the problem of learning forward and inverse mappings in the original spaces of velocity and waveforms can be reformulated as learning translations in their latent spaces. We provide theoretical justifications for the existence of these latent space translations in the Appendix A. Note that these assumptions can also be relaxed in specific instantiations of GFI. For example, the manifold assumption can be relaxed to learn latent spaces solely for translation and not reconstruction.
@@ 3.2 Prior Works in DL4SI as Special Cases of GFI
Table 1 summarizes how previous works in DL4SI can be viewed as specific realizations of the unifying GFI Framework, as described in the following. Schematic illustrations of all prior works in the framework of GFI are provided in the Appendix Figure 9.
@@
InversionNet and VelocityGAN: Both InversionNet and VelocityGAN use the same encoder-decoder architecture for learning the inverse mapping from seismic waveforms to velocity maps. They can be viewed as using an encoder $\mathtt{E}_{p}$ to map waveforms to a lower-dimensional space (represented as a flat vector), which is decoded back to velocity using decoder $\mathtt{D}_{v}$ (both encoder and decoder are implemented as ConvNets). The latent space translation is effectively the identity mapping, $L_{\tilde{p}\rightarrow\tilde{v}}=I$. Also, since they operate in inverse-only mode, they do not involve $\mathtt{E}_{v}$, $\mathtt{D}_{p}$, and $L_{\tilde{v}\rightarrow\tilde{p}}$ components in their architecture.
As a result, they do not perform manifold learning since the latent spaces of $\tilde{p}$ and $\tilde{v}$ are solely used for translation and not for reconstruction.
Auto-Linear: Auto-Linear employs a two-stage approach for learning both forward and inverse mappings using two disjoint latent space translations. In the first stage, self-supervised learning using reconstruction loss is utilized to train encoder-decoder pairs independently on velocity maps (i.e., $\mathtt{E}_{v}$ and $\mathtt{D}_{v}$) and seismic waveforms (i.e., $\mathtt{E}_{p}$ and $\mathtt{D}_{p}$). Note that the sizes of $\tilde{v}$ and $\tilde{p}$ are kept different from each other and higher than that of $v$ and $p$. In the second stage, two linear transforms are separately trained for the forward and inverse translations in the latent space, $L_{\tilde{v}\rightarrow\tilde{p}}$ and $L_{\tilde{p}\rightarrow\tilde{v}}$, while keeping the previously trained encoders and decoders frozen. Auto-Linear leverages the manifold assumption as the latent manifolds $\tilde{v}$ and $\tilde{p}$ are useful for reconstruction.
@@ 3.3 Proposed Model: Latent U-Net
We propose a novel architecture for solving forward and inverse problems using two latent space translation models implemented using U-Nets, termed Latent U-Net. As shown in Figure 2(a), Latent U-Net uses ConvNet backbones for both encoder-decoder pairs: ($\mathtt{E}_{v}$, $\mathtt{D}_{v}$) and ($\mathtt{E}_{p}$, $\mathtt{D}_{p}$), to project $v$ and $p$ to lower-dimensional representations. We also constrain the sizes of the latent spaces of $\tilde{v}$ and $\tilde{p}$ to be identical, i.e., $\text{dim}(\tilde{v})=\text{dim}(\tilde{p})$, so that we can train two separate U-Net models to implement the latent space mappings $L_{\tilde{v}\rightarrow\tilde{p}}$ and $L_{\tilde{p}\rightarrow\tilde{v}}$. Note that while U-Nets cannot be applied directly in the original spaces of velocity and waveforms due to a mismatch in dimensionalities, Latent U-Net allows us to bring to bear the power of U-Net in the domain of subsurface imaging, which has been very successful in several applications of computer vision.
We train Latent U-Net in an end-to-end manner to directly learn latent spaces useful for translation. This is in contrast to the two-stage training process employed in Auto-Linear where latent spaces are first pre-trained for reconstruction, followed by translation. Specifically, for the forward problem, the components $\mathtt{D}_{p}$, $L_{\tilde{v}\rightarrow\tilde{p}}$, and $\mathtt{E}_{v}$ are trained together by optimizing Equation 4, while for the inverse problem, the components $\mathtt{D}_{v}$, $L_{\tilde{p}\rightarrow\tilde{v}}$, and $\mathtt{E}_{p}$ are trained together by optimizing Equation 5. Since the training of forward and inverse modeling components of Latent U-Net are disjoint from each other, Latent U-Net does not explicitly learns manifolds useful for reconstruction. However, it is possible to modify the training objective of Latent U-Net by adding reconstruction losses in both forward and inverse problems to perform manifold learning.
@@ 3.4 Proposed Model: Invertible X-Net
Note that Invertible X-Nets facilitate bidirectional training as both the forward and inverse modeling components are trained simultaneously. As a result, both encoder-decoder pairs of Invertible X-Nets implicitly learn manifolds useful for reconstruction. Also, Invertible X-Net offers several key advantages over baselines. First, it simultaneously addresses both the forward and inverse problems within a single model architecture, whereas other baselines typically require training separate models for each task (e.g., Latent U-Net and Auto-Linear), leading to greater parameter efficiency. Additionally, the use of IU-Net ensures that the mappings between the latent spaces of velocity maps and seismic waveforms are bijective, guaranteeing a one-to-one mapping between these representations – a property not necessarily true for other models such as Latent U-nets and Auto-Linear. Further, the bidirectional training of the forward and inverse problems introduces a strong regularization effect as the gradients of the forward loss $\mathcal{L}_{\text{forward}}$ affects the parameters of $f_{\text{IU-Net}}$, thereby affecting both forward and inverse performance. Similarly, the gradients of the inverse loss $\mathcal{L}_{\text{inverse}}$ also impact both forward and inverse performance by updating the parameters of $f_{\text{IU-Net}}$.
@@ 4.1 Datasets
We consider the OpenFWI collection of datasets (Deng et al., 2022), which encompasses multi-structural benchmark datasets for DL4SI organized into three groups: Vel Family, Fault Family, and Style Family. The Vel Family consists of four datasets representing simple geological patterns, while the Fault Family includes four datasets with fault-like deformations, presenting greater modeling challenges. The Style Family utilizes a style transfer method to create a range of highly complex geological settings. The waveform data is structured as (# source $\times$ recording time $\times$ receiver length), resulting in a shape of ($5\times 1000\times 70$) for five uniformly spaced seismic sources and 70 receivers over 1000 milliseconds. In contrast, the velocity maps are represented as (1 $\times$ depth $\times$ receiver length), with a shape of ($1\times 70\times 70$) reflecting spatial dimensions of depth and horizontal coverage. We also conduct out-of-distribution evaluations on the Marmousi (Martin et al., 2006) and Overthrust datasets (Aminzadeh et al., 1996). They are standard benchmarks that closely mimic real-world scenarios, significantly surpassing the complexity of the OpenFWI datasets.
@@ 4.2 Model Architectures
Latent U-Net and Invertible X-Net: We use two variants for Latent U-Net: (1) (Small) with 17M parameters, and (2) (Large) with 33M parameters. For Invertible X-Net, we used an architecture with 24M parameters trained with and without cycle loss. See Appendix for details.
Baselines: For the inverse problem, we considered InversionNet, VelocityGAN, and Auto-Linear as baselines. We obtained their trained model checkpoints to reproduce their results. For the forward problem, we implemented FNO (Li et al., 2020) and a U-Net model (termed WaveformNet) as baseline models (see Appendix for details), along with using Auto-Linear.
Training Details: The input seismic waveforms were standardized using a StandardScaler to center the data with a zero mean and unit standard deviation, while the velocity data were scaled through MinMax normalization. See Appendix for details on hyper-parameter settings. To make a fair comparison with baseline methods, we kept the original train-test splits and reported results in the unnormalized space of both velocity and waveforms. We used Mean Absolute Error (MAE), Mean Square Error (MSE), and Structured Similarity (SSIM) as evaluation metrics.

### v3

@@ Abstract
In subsurface imaging, learning the mapping from velocity maps to seismic waveforms (forward problem) and waveforms to velocity (inverse problem) is important for several applications. While traditional techniques for solving forward and inverse problems are computationally prohibitive, there is a growing interest in leveraging recent advances in deep learning to learn the mapping between velocity maps and seismic waveform images directly from data. Despite the variety of architectures explored in previous works, several open questions remain unanswered such as the effect of latent space sizes, the importance of manifold learning, the complexity of translation models, and the value of jointly solving forward and inverse problems. We propose a unified framework to systematically characterize prior research in this area termed the Generalized Forward-Inverse (GFI) framework, building on the assumption of manifolds and latent space translations. We show that GFI encompasses previous works in deep learning for subsurface imaging, which can be viewed as specific instantiations of GFI. We also propose two new model architectures within the framework of GFI: Latent U-Net and Invertible X-Net, leveraging the power of U-Nets for domain translation and the ability of IU-Nets to simultaneously learn forward and inverse translations, respectively. We show that our proposed models achieve state-of-the-art performance for forward and inverse problems on a wide range of synthetic datasets and also investigate their zero-shot effectiveness on two real-world-like datasets. The code is available at https://github.com/KGML-lab/Generalized-Forward-Inverse-Framework-for-DL4SI
@@ 1 Introduction
The goal of subsurface imaging is to identify the structure and geophysical properties of layers underneath the Earth’s surface such as velocity maps ($v$), which is useful for a number of applications including energy exploration, carbon capture and sequestration, and developing earthquake early warning systems (Zhang et al., 2013). A typical approach for subsurface imaging is to conduct seismic surveys, where the elastic wave energy generated from one or more controlled sources is made to propagate through the Earth’s layers that are refracted, reflected, and received as seismic waveforms ($p$) on the surface by multiple receivers. Mathematically, the relation between velocity maps and seismic waveforms is governed by the following acoustic wave equation:
@@
For example, several deep learning methods have recently been proposed for FWI using encoder-decoder architectures such as InversionNet (Wu & Lin, 2019) and VelocityGAN (Zhang et al., 2019), where the encoder maps seismic data to latent feature representations that are decoded back to the velocity domain (Lin et al., 2023). Techniques for solving the forward problem include the seminal work on physics-informed neural networks (PINNs) (Raissi et al., 2019) and recent works such as Fourier Neural Operators (Li et al., 2020). A recent work has also pursued solving both forward and inverse problems together using a common architecture, Auto-Linear (Feng et al., 2024), where two auto-encoders are first separately trained to project velocity and waveform fields to higher-dimensional manifolds, which are then linked via linear transforms for manifold translation.
@@ 2.1 Deep Learning for Subsurface Imaging (DL4SI)
Since ground-truth observations of velocity maps are difficult to obtain in real-world settings, a number of unsupervised and self-supervised formulations have also been explored for FWI. This includes the framework of UPFWI (Jin et al., 2021) that uses the supervision of a physics-based forward model of the wave equation to enforce cycle-consistency with a data-driven inverse model. While the training of UPFWI is completely unsupervised and achieves comparable performance as InversionNet in some settings, it requires computationally expensive simulations of a forward model at every training epoch, restricting its scalability. Recent advancements, such as IFWI Du et al. (2024), FWIPLR Sun et al. (2023), and FWIRLG Yang & Ma (2023), integrate physics-based constraints and deep learning to tackle challenges like low-quality data and geological complexity. A recent line of work in semi-supervised learning for FWI involves the framework of Auto-Linear (Feng et al., 2024). In this framework, both velocity and waveforms are first mapped to higher-dimensional manifolds using two auto-encoders separately trained to minimize reconstruction loss in their respective domains. This is followed by a supervised learning of translations between the manifolds using two independent linear transforms (one for forward and one for inverse), building on previous observations of linearity in high-dimensional spaces for FWI (Feng et al., 2022).
@@ 3.1 Generalized Forward Inverse Framework
As shown schematically in Figure 1, our proposed framework of GFI unifies the learning of the forward problem $f_{v\rightarrow p}:\mathcal{V}\rightarrow\mathcal{P}$ and the inverse problem $f_{p\rightarrow v}:\mathcal{P}\rightarrow\mathcal{V}$ using a common architecture involving latent space translations. The GFI framework is inspired by two key assumptions. First, according to the manifold assumption, we assume that the velocity maps $v\in\mathcal{V}$ and seismic waveforms $p\in\mathcal{P}$ can be projected to their corresponding latent space representations, $\tilde{v}$ and $\tilde{p}$, respectively, which can be mapped back to their reconstructions in the original space, $\hat{v}$ and $\hat{p}$. Note that the sizes of the latent spaces can be smaller or larger than the original spaces. Further, the size of $\tilde{v}$ may not match with the size of $\tilde{p}$ (e.g., in Auto-Linear). Second, according to the latent space translation assumption, we assume that the problem of learning forward and inverse mappings in the original spaces of velocity and waveforms can be reformulated as learning translations in their latent spaces. We provide theoretical justifications for the existence of these latent space translations in the Appendix A. Note that these assumptions can also be relaxed in specific instantiations of GFI.
@@ 3.2 Prior Works in DL4SI as Special Cases of GFI
In this section, we revisit previous works in DL4SI and describe them using GFI terminologies, i.e., modeling mode (forward/inverse), size of latent space, whether they use manifold learning or not, and the nature of latent space translations. Table 1 summarizes prior works as specific realizations of the unifying GFI Framework. Schematic illustrations of all prior works in the framework of GFI are provided in the Appendix Figure 9.
@@
InversionNet and VelocityGAN: Both InversionNet and VelocityGAN operate exclusively in the inverse mode, mapping seismic waveforms to velocity maps. They utilize the same encoder-decoder architecture, with the encoder $\mathtt{E}_{p}$ projecting seismic waveforms ($p$) to a lower-dimensional latent space and the decoder $\mathtt{D}_{v}$ reconstructing the velocity maps ($v$) from this representation. The sizes of the latent spaces are low-dimensional, determined by the bottleneck layer of the U-Net architecture. In terms of latent space translation, these models rely on an identity mapping in the bottleneck layer, $L_{\tilde{p}\rightarrow\tilde{v}}=I$. However, we can also interpret them as having a U-Net for translation, if we view the first few and last few layers as the encoder-decoder pair while the intermediate layers perform translation. The sizes of the latent spaces are low-dimensional and they do not perform manifold learning since the latent spaces of $\tilde{p}$ and $\tilde{v}$ are solely used for translation and not for reconstruction.
Auto-Linear: Auto-Linear operates in both forward and inverse modes, using two disjoint latent space translations. It follows a two-stage approach: first, encoder-decoder pairs ( $\mathtt{E}_{v}$, $\mathtt{D}_{v}$, $\mathtt{E}_{p}$ and $\mathtt{D}_{p}$) are independently trained using reconstruction loss to project $v$ and $p$ into high-dimensional latent spaces ($\tilde{v}$ and $\tilde{p}$), where $dim(\tilde{v})\neq dim(\tilde{p})$. Second, linear transforms $L_{\tilde{v}\rightarrow\tilde{p}}$ and $L_{\tilde{p}\rightarrow\tilde{v}}$ are trained for translation, keeping the encoders and decoders frozen. Auto-Linear leverages the manifold assumption as the latent manifolds $\tilde{v}$ and $\tilde{p}$ are useful for reconstruction. The sizes of the latent spaces are higher-dimensional, designed specifically to facilitate linear mappings.
@@ 3.3 Proposed Model: Latent U-Net
We propose a novel architecture to solve forward and inverse problems using two latent space translation models implemented using U-Nets, termed Latent U-Net. As shown in Figure 2(a), Latent U-Net uses ConvNet backbones for both encoder-decoder pairs: ($\mathtt{E}_{v}$, $\mathtt{D}_{v}$) and ($\mathtt{E}_{p}$, $\mathtt{D}_{p}$), to project $v$ and $p$ to lower-dimensional representations. We also constrain the sizes of the latent spaces of $\tilde{v}$ and $\tilde{p}$ to be identical, i.e., $\text{dim}(\tilde{v})=\text{dim}(\tilde{p})$, so that we can train two separate U-Net models to implement the latent space mappings $L_{\tilde{v}\rightarrow\tilde{p}}$ and $L_{\tilde{p}\rightarrow\tilde{v}}$. Note that InversionNet can also be viewed as a special case of Latent U-Net in inverse-only mode with subtle architectural differences such as absence of skip connections, design of encoder/decoder and translation layers, and different dimensions of $\tilde{v}$ and $\tilde{p}$.
We train Latent U-Net in an end-to-end manner to directly learn latent spaces useful for translation. This is in contrast to the two-stage training process employed in Auto-Linear where latent spaces are first pre-trained for reconstruction, followed by translation. Specifically, for the forward problem, the components $\mathtt{D}_{p}$, $L_{\tilde{v}\rightarrow\tilde{p}}$, and $\mathtt{E}_{v}$ are trained together by optimizing Equation 4, while for the inverse problem, the components $\mathtt{D}_{v}$, $L_{\tilde{p}\rightarrow\tilde{v}}$, and $\mathtt{E}_{p}$ are trained together by optimizing Equation 5. Since the training of forward and inverse modeling components of Latent U-Net are disjoint from each other, Latent U-Net does not explicitly learns manifolds useful for reconstruction. However, it is possible to modify the training objective of Latent U-Net by adding reconstruction losses in both forward and inverse problems to perform manifold learning. See Appendix sections D.3.1 for details
@@ 3.4 Proposed Model: Invertible X-Net
Note that Invertible X-Nets facilitate bidirectional training as both the forward and inverse modeling components are trained simultaneously. As a result, both encoder-decoder pairs of Invertible X-Nets implicitly learn manifolds useful for reconstruction. Also, Invertible X-Net offers several key advantages over baselines. First, it simultaneously addresses both the forward and inverse problems within a single model architecture, whereas other baselines typically require training separate models for each task (e.g., Latent U-Net and Auto-Linear), leading to greater parameter efficiency. Additionally, the use of IU-Net ensures that the mappings between the latent spaces of velocity maps and seismic waveforms are bijective, guaranteeing a one-to-one mapping between these representations – a property not necessarily true for other models such as Latent U-nets and Auto-Linear. Further, the bidirectional training of the forward and inverse problems introduces a strong regularization effect as the gradients of the forward loss $\mathcal{L}_{\text{forward}}$ affects the parameters of $f_{\text{IU-Net}}$, thereby affecting both forward and inverse performance. Similarly, the gradients of the inverse loss $\mathcal{L}_{\text{inverse}}$ also impact both forward and inverse performance by updating the parameters of $f_{\text{IU-Net}}$. Further details about the Invertible X-Net architecture are provided in Appendix Section D.3.2.
@@ 4.1 Datasets
We consider the OpenFWI collection of datasets (Deng et al., 2022), comprising multi-structural benchmark datasets for DL4SI grouped into: Vel, Fault, and Style Families. The Vel Family includes four datasets with simple geological patterns, while the Fault Family contains four datasets with fault-like deformations, presenting greater modeling challenges. The Style Family uses style transfer methods to create complex geological settings. The waveform data is structured as (# source $\times$ recording time $\times$ receiver length), resulting in a shape of ($5\times 1000\times 70$) for five uniformly spaced seismic sources and 70 receivers over 1000 milliseconds. In contrast, the velocity maps are represented as (1 $\times$ depth $\times$ receiver length), with a shape of ($1\times 70\times 70$) reflecting spatial dimensions of depth and horizontal coverage. We also conduct out-of-distribution evaluations on the Marmousi (Martin et al., 2006) and Overthrust datasets (Aminzadeh et al., 1996). They are standard benchmarks that closely mimic real-world scenarios and exceed the complexity of OpenFWI datasets.
@@ 4.2 Model Architectures
Latent U-Net and Invertible X-Net: We use two variants for Latent U-Net: (1) (Small) with 17M parameters, and (2) (Large) with 33M parameters. For Invertible X-Net, we used an architecture with 24M parameters trained with and without cycle loss. See Appendix D.1 and D.3for details. Baselines: For the inverse problem, we considered InversionNet, VelocityGAN, and Auto-Linear as baselines. We obtained their trained model checkpoints to reproduce their results. For the forward problem, we implemented FNO (Li et al., 2020) and a U-Net model (termed WaveformNet) as baseline models (see Appendix C for details), along with using Auto-Linear.
Training Details: The input seismic waveforms were standardized using a StandardScaler to center the data with a zero mean and unit standard deviation, while the velocity data were scaled through MinMax normalization. See Appendix D.2 for details on hyper-parameter settings. To make a fair comparison with baseline methods, we kept the original train-test splits and reported results in the unnormalized space of both velocity and waveforms. We used Mean Absolute Error (MAE), Mean Square Error (MSE), and Structured Similarity (SSIM) as evaluation metrics since neither metric alone is fully comprehensive. MAE captures pixel-level accuracy while SSIM highlights structural similarity.